The number of entries made in accounts receivable is a poisson distribution with average 2 per day .What is the probability that on a specified day at least 3 entries will be received?     


1
Expert's answer
2021-05-27T09:40:20-0400

Given,

"Mean\ \mu =2"

and we know that,

"For\ Poisson\ Distribution\\P(x)=\dfrac{\mu^xe^{-\mu}}{x!}\\Putting\ \mu=2\\P(x)=\dfrac{2^xe^{-2}}{x!}"


Now, Probability that on a specified day at least 3 entries will receive = "P(x\geq 3)"


"P(x\geq 3)=1-[P(0)+P(1)+P(2)]"

"=1-\dfrac{1}{e^2}[\dfrac{2^0}{0!}+\dfrac{2^1}{1!}+\dfrac{2^2}{2!}]\\\ \\=1-\dfrac{1}{e^2}[1+2+2]\\\ \\=1-\dfrac{5}{e^2}=0.323"


Hence, "\boxed{P(x\geq 3)=0.323}"